OpenStudy (anonymous):

Given the trinomial 2x2 + 4x − 2, what is the value of the discriminant?

OpenStudy (anonymous):

2x^2+4x-2

OpenStudy (anonymous):

For the trinomial \(ax^2+bx+c\) the discriminant is \(b^2-4ac\) Can you pick out the values for a, b, and c from your trinomial?

OpenStudy (anonymous):

4^2-4(2)(-2)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would be 32 right? 4^2=16 then -4*2=-8 -8*-2=16 then 16+16=32

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thank you! they didn't have that formula in the lesson made it easier.

OpenStudy (anonymous):

you're welcome